OpenStudy (lgbasallote):

i need algebra help....i have a formula \[\huge \frac{r_2}{r_1} = \frac{t_1}{t_2} = \sqrt{\frac{M_1}{M_2}}\] t_1 = 15 t_2 = 12 M_2 = 28.02 is it possible to get the value of r2?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i don't think so

OpenStudy (lgbasallote):

why not?

OpenStudy (lgbasallote):

what do i need to get r2?

OpenStudy (anonymous):

r1

OpenStudy (lgbasallote):

and i suppose i cant get r1 either huh

OpenStudy (lgbasallote):

what if the formula was \[\huge \frac{r_2}{r_1} = \frac{\frac{V_2}{t_2}}{\frac{V_1}{t_1}} = \sqrt{\frac{M_1}{M_2}}\] where V)2 = v_1 is it possible then?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

still... no

OpenStudy (lgbasallote):

wew...is it possible to get the value of V?

OpenStudy (anonymous):

i would think not

OpenStudy (lgbasallote):

ugh >.<

OpenStudy (anonymous):

lol whatever you have for the middle thing will only give you the propotion of the other items

OpenStudy (anonymous):

and btw why do you need this?

OpenStudy (lgbasallote):

because i need to find the root-mean-square speed of a compound...it's a chem question...but since im just looking for an algebraic way to isolate r2 i figured math is a better place to ask

OpenStudy (amistre64):

it looks like you essentially have: y = mx; and are asking to find the value for y given that you know the slope only

OpenStudy (lgbasallote):

uhhh is that a mispost @amistre64 ?

OpenStudy (amistre64):

more like a mister.post ....

OpenStudy (lgbasallote):

lol i'll take that as a yes

OpenStudy (amistre64):

\[y = mx\] \[\frac yx=m\] m=t1/t2 y/x = r2/r1

OpenStudy (lgbasallote):

anyway like i said this isnt really an algebra problem...it's a chemistry problem but i need to isolate r2

OpenStudy (amistre64):

\[r2 = \frac{t2}{t1}r1\]

OpenStudy (amistre64):

or however teh ts are stacked ...

OpenStudy (amistre64):

we could also compare this to the law of sines if we want

OpenStudy (lgbasallote):

okay...i think this is getting overcomplicated now o.O

OpenStudy (amistre64):

you can determine an M part; but not a distinct r part is all

OpenStudy (lgbasallote):

....so in layman's terms it's impossible?

OpenStudy (amistre64):

correct, it is impossible

OpenStudy (lgbasallote):

you couldve said that sooner :p lol